Please help fill in the highlighted values below:Weighted Average Cost of Capital | D/E Wd 0.3 Rd 0% 4.77% 5.25% 6.75% 0.231 0.333 Rd(1-1) b 0.00% 3.15% 0.03465 4.46% 1.24 1.
Answer #1

The highlighted columns can be filled as shown in below table.

Rd(1-T) Wce A 1 D/E 20 3 0.3 4 =B4/F4 5 0.75 B Wd JO 0.231 0.333 =A5*F5 C Rd 10 0.0477 0.0525 0.0675 1.24 1.48 1.64 =E2*(1+(1

The resultant values are mentioned in below table.

E b 1 2 3 4 5 A D/E 0 0.3 0.499 0.75 B Wd 0 0.231 0.333 0.428 C Rd 0% 4.77% 5.25% 6.75% D Rd(1-1) 0% 3.15% 3.47% 4.46% 1.24 1